OpenStudy (anonymous):

Where am I messing up? Use logarithmic differentiation to find the derivative of the function. y= sqrt(x)^7/x ------------------- I will post my work below.

OpenStudy (turingtest):

\[ y= \sqrt{x^7}/x=x^{7/2}/x=x^{5/2}\]if your formula is written correctly you can first simplify it this way. Is this written correctly?

OpenStudy (anonymous):

\[lny=lnx^{7x/2}\] \[u=7x/2\] \[lny=ulnx\] \[(1/y) (dy/dx) = (7x/2)(1/x)+(lnx)(7/2)\] \[(dy/dx) = [(7x/2)(1/x)+(lnx)(7/2)] [(\sqrt{x})^{7x}]\]

OpenStudy (turingtest):

so x is in the exponent too?

OpenStudy (anonymous):

Oops, the question is actually, i wrote it wrong. \[(\sqrt{x})^{7x}\]

OpenStudy (anonymous):

I'm being told that my final answer is wrong. Yet i can't figure out where I am messing up.

OpenStudy (turingtest):

\[y=\sqrt{x}^{7x}\] right? let me try it for a minute...

OpenStudy (anonymous):

Yes Turing, that's right. By the way, I really appreciate you taking the time to help me.

OpenStudy (turingtest):

I got the same as you but noticed that in the last line the (7x/2)(1/x) cancels to 7/2 so \[y'=(7/2)(1+\ln x)\sqrt{x}^{7x}\]I hope that works...

OpenStudy (anonymous):

Boom, that was the problem. Thanks, you're awesome!

OpenStudy (turingtest):

Sweet, but you did it pretty much yourself. I just saw a detail, that's why two heads are better than one :)